My sister is half my age.
Next year she will be three times as old as I was thirteen years ago.
How old am I now?

Answer by chibisan(131)

Your age = 2x , sister = x
1+x = 3(2x-13)
1+x = 6x - 39
40 = 5x
x = 8
2x = 16 (your age)